### v1

@@ 4.2 Jet emission model
The accelerated electrons and protons inside the jet are cooled down primarily via synchrotron emission, which covers the regime from radio to soft gamma-ray wavelengths. On the other hand, the hadronic emission mechanism involves collisions between the relativistic protons and thermal ones, which produce pions and muons that decay to detectable neutrinos and gamma-rays.
As described in the previous subsection, we consider an IC spectrum component to explain the MeV tail observed from Cygnus X-1. We employ the mathematical formulation described in Rybicki & Lightman (2008) for those calculations, assuming both the Compton and inverse Compton cases. We adopt the conclusions of Blumenthal & Gould (1970) to estimate the scattered spectrum caused by an electron distribution with $\gamma_{e}\gg 1$ while considering both the Thomson and the Klein-Nishina limits.
4.3 Hard/Low spectral state
@@
The synchrotron component due to highly relativistic electrons and protons (the latter achieve energies up to $E_{max}\approx 10^{7}$ GeV) confirms the Fermi-LAT gamma-ray measurements and upper limits, as seen in Fig. 11. At last, the proton-proton mechanism is too weak in our lepton-dominated jet model (i.e., $n_{p}/n_{e}=0.009$) to reach the upper limits set by the MAGIC collaboration for the high-energy gamma-ray regime (Ahnen et al., 2017). Nevertheless, photo-absorption caused by the disk is dominant near the lower-jet emitting regions. Hence, detected gamma-ray fluxes above 10 GeV are more likely to originate inside the expanded jet frame (Zanin et al., 2016).
@@
4.4 Deviation from the kerrbb spectra
There are three primary sources of error concerning the approach presented in this work for calculating relativistic spectra. The first is the intrinsic error margin of the analytic function $f(\theta,\alpha_{*})$, which is relatively small and does not exceed 1% (Campitiello et al., 2018). The second one is due to the PN-potential adoption, which can reach 8-9% in the case of the Artemova solution. Lastly, the error margin introduced in the context of this work is due to the deviation between the PN-generated $\eta_{rad}(R_{in})$ and Eq. (14), with the choice of the $p=1.17$ index.
[…]
@@ 5 Summary and conclusions
The energy distribution radiated away from a thin accretion disk described by the kerrbb model is equivalent to the one emitted by a Shakura-Suniaev disk with a modified innermost radius, which depends on the system’s spin and viewing angle. It is crucial to use the spin-dependent PN potential. Otherwise, the modified boundary cannot be applied in XRBs inclined by $\theta>60^{\circ}$, independently of the black hole’s rotation. The best-suited potential is the one suggested by Artemova et al. (1996) because it produces valid results for small radii, even more so than the $R_{ISCO}$. Hence, the boundary transition, unconstrained by the spin parameter, is within the $R_{in}/R_{ISCO}=0.2-1.0$ range, when $\theta\geq 60^{\circ}$, and $R_{in}/R_{ISCO}=1.0-2.0$, otherwise.
The thermal emission from Cygnus X-1 in the high/soft state is reproduced by a black-body spectrum associated with the inner disk surface. The non-thermal tail at higher energies is fitted by the IC scattered emission of the incident disk photons due to a thermal/non-thermal corona hybrid. Regarding the hard/low state, the synchrotron emission from a fairly simple lepto-hadronic jet is sufficient to match the radio and soft gamma-ray measurements. The soft X-ray emission approximates the black-body spectrum from a truncated accretion disk. It is accompanied by a harder IC component due to a more extensive, hotter, and optically thicker corona compared to the one in the soft state. Shifting to higher frequencies, the collisions of protons inside the jet are incapable of generating detectable high-energy gamma-rays.
Our results show a very good agreement with the fully relativistic treatment of the Novikov-Thorne model incorporated in the kerrbb code, with a deviation within 0.03-4%. The intrinsic error of the pseudo-potential, which leads to an overestimated radiated luminosity, is minimized by the consistent boundary transition to fit the relativistic predictions. However, the inaccuracy in the radiative efficiency, assuming an inner radius of $R_{ISCO}$, remains. Therefore, we propose a fixing factor of $k\approx 1.075$ that marginally shifts the $R_{ISCO}$ to match the theoretical predictions.

### v2

@@ 4.3 Jet emission model
The accelerated electrons and protons inside the jet are cooled down primarily via synchrotron emission, which covers the regime from radio to soft gamma-ray wavelengths. On the other hand, the hadronic emission mechanism involves collisions between the relativistic protons and thermal ones, which produce pions and muons that decay to high-energy neutrinos and gamma-rays.
As described in the previous subsection, we consider an IC spectrum component to explain the MeV tail observed from Cygnus X-1. We employ the mathematical formalism described in Rybicki & Lightman (2008) for those calculations, assuming both the Compton and inverse Compton cases. We adopt the conclusions of Blumenthal & Gould (1970) to estimate the scattered spectrum caused by an electron distribution with $\gamma_{e}\gg 1$ while considering both the Thomson and the Klein-Nishina limits.
4.4 Cygnus X-1: hard/low spectral state
@@
The synchrotron component due to highly relativistic electrons and protons (the latter achieve energies up to $E_{max}\approx 10^{7}$ GeV) confirms the Fermi-LAT gamma-ray measurements and upper limits, as seen in Fig. 13. At last, the proton-proton mechanism is too weak in our lepton-dominated jet model (i.e., $L_{p}/L_{e}=0.009$) to reach the upper limits set by the MAGIC collaboration for the high-energy gamma-ray regime (Ahnen et al., 2017). Nevertheless, photo-absorption caused by the disk is dominant near the lower-jet emitting regions. Hence, detected gamma-ray fluxes above 10 GeV are more likely to originate inside the expanded jet frame (Zanin et al., 2016; Papavasileiou et al., 2023).
@@ 5 Summary and conclusions
The energy distribution radiated away from a thin accretion disk described by the Novikov-Thorne and kerrbb models is equivalent to the one emitted by a Shakura-Sunyaev disk with a modified innermost radius, which depends on the system’s spin and viewing angle. It is crucial to use the spin-dependent PN potential. Otherwise, the modified boundary cannot be applied in XRBs inclined by $\theta>60^{\circ}$, independently of the black hole’s rotation. The best-suited potential is the one suggested by Artemova et al. (1996) because it produces valid results for small radii, even more so than the $R_{ISCO}$. Hence, the boundary transition, unconstrained by the spin parameter, is within the $R_{in}/R_{ISCO}=0.2-1.0$ range, when $\theta\geq 60^{\circ}$, and $R_{in}/R_{ISCO}=1.0-2.0$, otherwise.
The thermal emission from Cygnus X-1 in the high/soft state is reproduced by a black-body spectrum associated with the inner disk surface. The non-thermal tail at higher energies is fitted by the IC scattered emission of the incident disk photons due to a thermal/non-thermal corona hybrid. Regarding the hard/low state, the synchrotron emission from a fairly simple lepto-hadronic jet is sufficient to match the radio and soft gamma-ray measurements. The soft X-ray emission approximates the black-body spectrum from a truncated accretion disk. It is accompanied by a harder IC component due to a more extensive, hotter, and optically thicker corona compared to the one in the soft state. Regarding higher energies, the collisions of protons inside the jet are incapable of generating detectable high-energy gamma-rays.
Our results show a very good agreement with the fully relativistic treatment of the Novikov-Thorne model incorporated in the kerrbb code, with a deviation within 0.03-4%. The intrinsic error of the pseudo-potential, which leads to an overestimation of the radiated luminosity, is minimized by the consistent boundary transition to fit the relativistic predictions. However, the inaccuracy in the radiative efficiency, assuming an inner radius of $R_{ISCO}$, remains. Therefore, we propose a fixing factor of $k\approx 1.075$ that marginally shifts the $R_{ISCO}$ to match the theoretical predictions.